The Welcher Adult Intelligence Test Scale is composed of several subtexts. On one subtest the raw scores have a mean of 55 and a variance of 25.Assuming these raw scores are from a normal distribution.

a) what is the probability of obtaining a raw score between the mean and 46?

b)what is the probability of obtaining a raw score between 50 and 58?


Show all calculations!


1
Expert's answer
2021-05-24T19:01:31-0400

"\mu=55,\sigma^2=25"


Standard deviation "=\sqrt{25}=5"


(a)  Probability of obtaining a raw score between the mean and 46-


z score for mean "z=\dfrac{55-55}{5}=0"


z-score for 46 "z=\dfrac{46-55}{5}=-1.8"


Probability is"=P(-1.8<z<0)=0.464"


(b) Probability of obtaining a raw score between 50 and 58-


"=P(50<X<58)\\[9pt]=P(\dfrac{50-55}{5}<z<\dfrac{58-55}{5})\\[9pt]=P(-1<z<0.6)\\[9pt]=0.567"